# Factor Models of Matrix-Valued Time Series: Nonstationarity and Cointegration

### v1

@@ 5.2 An empirical application
In this section, we apply the proposed model and methodology to analyze international trade flows. By modeling possibly nonstationary trade volumes directly, we aim to examine the evolution pattern of global trade and the international trade network, and assess the impact of major events such as the global financial crisis and the COVID-19 pandemic. Our empirical analysis is based on monthly bilateral export volumes of commodity goods among 23 countries and regions, obtained from the International Monetary Fund’s Direction of Trade Statistics. The data cover free-on-board export values denominated in US dollars over a 273-month period from January 2000 to September 2022. The raw trade volume data are log-transformed, whereby the first differences correspond to the growth rates of international trade flows. The countries and regions include Australia (AUS), Canada (CAN), Mainland China (CHN), Denmark (DNK), Finland (FIN), France (FRA), Germany (DEU), Hong Kong China (HKG), Indonesia (IDN), Ireland (IRL), Italy (ITA), Japan (JPN), Korea (KOR), Malaysia (MYS), Mexico (MEX), Netherlands (NLD), New Zealand (NZL), Singapore (SGP), Spain (ESP), Sweden (SWE), Thailand (THA), United Kingdom (GBR), and United States (USA).
@@
Figure 2 plots the time series of the estimated common factor. From a modeling perspective, the strong trending pattern in the factor reflect its nonstationary feature. Model (2.2) explicitly accommodates integrated components, thereby capturing persistent global movements in international trade. In this context, the estimated factor can be interpreted as a global demand index: bilateral trade volumes fall worldwide when it drops; and international trade expands globally when it rises. Thus, the model effectively identifies the dominant source of co-movement in international trade flows, highlighting its relevance for capturing global shocks in a parsimonious and interpretable manner. Furthermore, the dynamic behavior of the estimated factor aligns closely with several major global events. The plot shows a persistent upward trend beginning in 2002, which may be attributed to China’s accession to the World Trade Organization and the subsequent acceleration in its trade activities with major economies (e.g., Hsieh and Ossa, 2016). Notably, it exhibits a sharp drop around 2009, corresponding to the so-called “Great Trade Collapse” due to the 2008 Global Financial Crisis. The estimated factor also drops sharply in early 2020 due to the outbreak of COVID-19 pandemic, followed by a quick rebound, reflecting the V-shaped pattern of global trade disruption and recovery. In addition, a smaller but visible decline is observed around 2015–2016, coinciding with the documented slowdown in global trade during that period (e.g., Constantinescu et al, 2020).
@@ 6 Conclusion
We have introduced a general matrix factor model to tackle large-scale trending matrix time series, making full case of the intrinsic matrix data structure and aiming to improve the estimation convergence. This is substantially different from the conventional vectorized factor model framework which transforms the matrix time series observations into vectors and often results in loss of sample information. We propose two types of PCA-based estimation techniques: mPCA and mPANIC, according to different nonstationary features in matrix common and idiosyncratic components. Under various model assumptions, we establish the convergence theory for the estimated factor loading matrices and nonstationary factor matrices and the consistency of the factor number estimation via the eigenvalue ratio criterion. In particular, we allow the existence of weak factors in the developed factor model structure, where the factor strength can be heterogeneous. The obtained mean squared convergence rates (for factor loading matrix estimation) and uniform convergence rates (for factor matrix estimation) depend on the matrix size $(p_{1},p_{2})$, time series length $T$ and the weak factor strength. The Monte-Carlo simulation study justifies the estimation convergence in finite samples. The empirical application to international trade flows shows that the estimated factor can be viewed as the global trade index and the estimated row and column loading matrices are useful to describe bilateral trade linkages between major economies (such as China and the United States).
@@
In this appendix, we provide the detailed proofs of the main theoretical results. Throughout the proofs, we let $C$ denote a generic positive constant whose value may change from line to line.
@@
We next give some technical lemmas together with their proofs. These lemmas play a crucial role in proving the main theoretical results in Appendix A.
@@
Suppose that Assumptions 3.1(ii), 3.53.7 are satisfied. Then we have
@@
Suppose that Assumptions 3.1(ii), 3.53.7 are satisfied. The rotation matrices $\overline{{\boldsymbol{H}}}_{R}$ and $\overline{{\boldsymbol{H}}}_{C}$ are invertible w.p.a.1, and have the following convergence properties:
@@
Proof of Lemma B.14.  To save the space, we prove the first assertion in (B.38). It follows from (A.2) that

### v2

@@ 5.2 An empirical application
In this section, we apply the proposed model and methodology to analyze international trade flows. By modeling possibly nonstationary trade volumes directly, we aim to examine the evolution pattern of global trade and the international trade network, and further assess the impact of major events such as the global financial crisis and the COVID-19 pandemic. Our empirical analysis is based on monthly bilateral export volumes of commodity goods among 23 countries and regions, obtained from the International Monetary Fund’s Direction of Trade Statistics. The data cover free-on-board export values denominated in US dollars over a 273-month period from January 2000 to September 2022. The raw trade volume data are log-transformed, whereby the first differences correspond to the growth rates of international trade flows. The countries and regions include Australia (AUS), Canada (CAN), Mainland China (CHN), Denmark (DNK), Finland (FIN), France (FRA), Germany (DEU), Hong Kong China (HKG), Indonesia (IDN), Ireland (IRL), Italy (ITA), Japan (JPN), Korea (KOR), Malaysia (MYS), Mexico (MEX), Netherlands (NLD), New Zealand (NZL), Singapore (SGP), Spain (ESP), Sweden (SWE), Thailand (THA), United Kingdom (GBR), and United States (USA).
@@
Figure 2 plots the time series of the estimated common factor. From a modeling perspective, the strong trending pattern in the factor reflects its nonstationary feature. Model (2.2) explicitly accommodates integrated components, thereby capturing persistent global movements in international trade. In this context, the estimated factor can be interpreted as a global demand index: bilateral trade volumes fall worldwide when it drops; and international trade expands globally when it rises. Thus, the model effectively identifies the dominant source of co-movement in international trade flows, highlighting its relevance for capturing global shocks in a parsimonious and interpretable manner. Furthermore, the dynamic behavior of the estimated factor aligns closely with several major global events. The plot shows a persistent upward trend beginning in 2002, which may be attributed to China’s accession to the World Trade Organization and the subsequent acceleration in its trade activities with major economies (e.g., Hsieh and Ossa, 2016). Notably, it exhibits a sharp drop around 2009, corresponding to the so-called “Great Trade Collapse” due to the 2008 Global Financial Crisis. The estimated factor also drops sharply in early 2020 due to the outbreak of COVID-19 pandemic, followed by a quick rebound, reflecting the V-shaped pattern of global trade disruption and recovery. In addition, a smaller but visible decline is observed around 2015–2016, coinciding with the documented slowdown in global trade during that period (e.g., Constantinescu et al, 2020).
@@ 6 Conclusion
We have introduced a general matrix factor model to tackle large-scale trending matrix time series, making full case of the intrinsic matrix data structure and aiming to improve the estimation convergence. This is substantially different from the conventional vectorized factor model framework which transforms the matrix time series observations into vectors and often results in loss of sample information. We propose two types of PCA-based estimation techniques: mPCA and mPANIC, according to different nonstationary features in matrix common and idiosyncratic components. Under various model assumptions, we establish the convergence theory for the estimated factor loading matrices and nonstationary factor matrices and the consistency of the factor number estimation via the eigenvalue ratio criterion. In particular, we allow the existence of weak factors in the proposed factor model structure, where the factor strength can be heterogeneous. The obtained mean squared convergence rates (for factor loading matrix estimation) and uniform convergence rates (for factor matrix estimation) depend on the matrix size $(p_{1},p_{2})$, time series length $T$ and the weak factor strength. The Monte-Carlo simulation study justifies the estimation convergence in finite samples. The empirical application to international trade flows shows that the estimated factor can be viewed as the global trade index and the estimated row and column loading matrices are useful to describe bilateral trade linkages between major economies (such as China and the United States).
Supplement
The supplemental document contains proofs of the main theorems (in Appendix A) and some technical lemmas with proofs (in Appendix B).
Supplement to “Factor Models of Matrix-Valued Time Series:
Nonstationarity and Cointegration”
Degui Li, Yayi Yan, Qiwei Yao
University of Macau, Shanghai University of Finance and Economics, London School of Economics
This supplemental file contains two appendices. Specifically, we provide the detailed proofs of the main theoretical results in Appendix A and present some technical lemmas together with their proofs in Appendix B. Throughout the proofs, we let $C$ denote a generic positive constant whose value may change from line to line.
@@
Suppose that Assumptions 3.1(ii) and 3.53.7 are satisfied. Then we have
@@
Suppose that Assumptions 3.1(ii) and 3.53.7 are satisfied. The rotation matrices $\overline{{\boldsymbol{H}}}_{R}$ and $\overline{{\boldsymbol{H}}}_{C}$ are invertible w.p.a.1, and have the following convergence properties:
@@
Proof of Lemma B.14.  To save the space, we only prove the first assertion in (B.38). It follows from (A.2) that
@@
$\widehat{\boldsymbol{H}}_{C}^{\ast}$ $=$ $\left(\frac{1}{p_{1}^{\alpha_{R,1}}T^{2}}\sum_{t=1}^{T}{\boldsymbol{F}}_{t}^{{}^{\intercal}}{\boldsymbol{R}}^{{}^{\intercal}}{\boldsymbol{R}}{\boldsymbol{F}}_{t}\right)\left({\boldsymbol{C}}^{{}^{\intercal}}\widehat{\boldsymbol{C}}\mathbf{B}_{C}^{-1}\right)\left(\widehat{\boldsymbol{V}}_{C}^{\ast}\right)^{-1}.$